What is the pH of a solution prepared by adding 1.59 g of sodium nitrite to 195 mL of water? Ka of HNO2 is 4.5 x 10-4.
Answer #1

solution Given 1.59 g of NaNo dissolved in 195 mL od water and formy a solution W = 1.59 g weight at Nanoz Gram molecular wei- log 4.5 - log 104 Kad HNO2 = 4.58 104 pka - log ka LO -log 4.5 x 10 10 LO - log4.5-(-4) log LO lo lo = -0.65 +4 IN 3.35 NaNpH = ² ( 14+3.35 - 11 - 0.92) - Ź (17.75 -0.92) = 16.83 2 - 8.415 pH d Nanoa solution = 8.415